OpenStudy (anonymous):

(ALGEBRA 1) A ball is thrown upward from the top of a building. The function below shows the height of the ball above the ground, f(t), in feet, at different times (t) in seconds: f(t) = -16t2 + 32t + 90 The average rate of change of f(x) from x = 4 to x = 6 is

OpenStudy (anonymous):

like finding the slope of the line compute \[\frac{f(6)-f(4)}{6-4}\]

OpenStudy (anonymous):

clear how to compute \(f(6)\) etc?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

k then lets go slow

OpenStudy (anonymous):

\[f(t) = -16t^2 + 32t + 90\]\[f(6)=-16(6^2)+32\times 6+90\]

OpenStudy (anonymous):

i would use a calculator lets try it

OpenStudy (anonymous):

wait a sec i get a negative number do you?

OpenStudy (anonymous):

im so confused.

OpenStudy (anonymous):

me too when you evaluate the function at \(t=6\) you seem to get a negative number normally that is not a problem, but in this equation \(f(t)\) is supposed to be the height, which cannot be negative

OpenStudy (anonymous):

i would say that the question makes no sense, because when the ball hits the ground, it stops

OpenStudy (anonymous):

So what do you think the answer is?

OpenStudy (anonymous):

we can write an answer, but it will make no sense but if you need an answer we can find one

OpenStudy (anonymous):

sorry i meant \[f(6)=-16(6^2)+32\times 6+90=-294\]

OpenStudy (anonymous):

\[f(4)=-16(4)^2+32\times 4+90=-38\]

OpenStudy (anonymous):

\[f(6)-f(4)=-294+38=-256\]

OpenStudy (anonymous):

\[\frac{f(6)-f(4)}{6-4}=\frac{-256}{2}=-128\]

OpenStudy (anonymous):

that is the answer it makes no sense, but that is the answer

OpenStudy (anonymous):

the real answer is that at \(3.6\) seconds the ball hits the ground, and goes no lower

OpenStudy (anonymous):

So the rate of change from x=4 to x=6 would be what

OpenStudy (anonymous):

@satellite73